Calculate the concentration of Fe2+ in each of the standard solutions listed below.

Standard Solution:

Standard # Volume of 2.50x10-4 M (NH4)2Fe(SO4)2 (mL) Volume of 0.0055 M 1,10 phenanthroline (mL)
1 2 5
2 4 5
3 6 5
4 9 5

Thanks!
Answer #1

Standard # 1

Total volume = 2 + 5 = 7 mL

[Fe^{2+}] = 2.50 \times 10^{-4} \: M \times \frac {2 \: mL}{7 \: mL} = 7.14 \times 10^{-5} \: M

Standard # 2

Total volume = 4 + 5 = 9 mL

[Fe^{2+}] = 2.50 \times 10^{-4} \: M \times \frac {4 \: mL}{9 \: mL} = 1.11 \times 10^{-4} \: M

Standard # 3

Total volume = 6 + 5 = 11 mL

[Fe^{2+}] = 2.50 \times 10^{-4} \: M \times \frac {6 \: mL}{11 \: mL} = 1.36 \times 10^{-4} \: M

Standard # 4

Total volume = 9 + 5 = 14 mL

[Fe^{2+}] = 2.50 \times 10^{-4} \: M \times \frac {9 \: mL}{14 \: mL} = 1.61 \times 10^{-4} \: M